Circular No. 8957

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2671 OPHIUCHI R. J. Rudy, D. K. Lynch, R. W. Russell, and A. M. Gilbert, The

Aerospace Corporation; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V2671 Oph on June 14 UT, obtained as above. V2671 Oph (Nova Oph 2008 No. 2; cf. IAUC 8951) is an “Fe II”- type nova with comparatively narrow lines of FWHM = 1200 km/s. It is at a very low stage of excitation, showing exceptionally bright lines of neutral carbon and strong emissison from the Ca II infrared triplet. The O I lines are also strong and indicate a large interstellar reddening of E(B − V) = 2.0. Weak, first-overtone emission from carbon monoxide beginning at 2.29 μm is present. No emission from dust is yet evident in the spectrum, but the presence of molecular gas suggests that its formation is imminent.

COMET C/2007 W1 (BOATTINI) A. J. Lovell, Agnes Scott College; and E. S. Howell, Arecibo Obser-

vatory, report detection of 18-cm OH lines in comet C/2007 W1 with the 100-m Green Bank Telescope. Hyperfine ratios of lines at 1667, 1665, 1612, and 1720 MHz are 8.6:4.8:1.1:1. Production rates log Q(OH) are as follows: June 13.75 UT, 28.20; 14.79, 28.21; 15.78, 28.27. Mapping at 8′ scale suggests that collisional quenching is present in the inner 20000 km of the coma. Line shapes are consistent with parent (water) outflow velocity of 0.78 km/s.

Visual total-magnitude estimates (cf. IAUC 8951) by R. Salvo, Montevideo, Uruguay (7×50 binoculars): June 8.95 UT, 5.6; 10.93, 5.5; 12.91, 5.5; 30.35, 6.0.

COMETS C/2008 G5 AND C/2008 G6 (SOHO) Further to IAUC 8956, additional apparent comets have been found on

SOHO website images. C/2008 G5 (a Kreutz sungrazer) was very diffuse, very faint (mag 8), and elongated. C/2008 G6, which was small and stellar in appearance (mag ∼ 7.5), was suggested by R. Kracht to be a return of the Kracht-group presumed-comet C/2002 S11 (cf. IAUC 7991), and an orbit by B. G. Marsden on MPEC 2008-L29 links the two apparitions with an orbital period of ∼ 5.54 yr.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 G5 Apr. 10.018 1^{h}24.^{m}4 + 7^{o}12′ C2 MU 2008-L29
C/2008 G6  12.921 1 31.5 + 9 25 C2 BZ 2008-L29
2008 July 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green